OpenStudy (anonymous):

In planet x the formula T=au gives the time of flight of a particle thrown vertically up with a velocity u in that planent the formula followed for final velocity of a particle dropped from rest from a height h is?

OpenStudy (anonymous):

What is \(T\)? What is \(a\)?

OpenStudy (anonymous):

if T is time a -acceleration and u is velocity then this formula is not correct, due to its dimensions. I'd write this formula in another way T = u/a

OpenStudy (anonymous):

@open_study1

OpenStudy (anonymous):

^Yup. So there are issues in this question's posing.

OpenStudy (anonymous):

it is said it take place in other planent

OpenStudy (anonymous):

Regardless of what planet this takes place in, unless it is outside of our universe and part of another universe in the multiverse, the laws of Physics still apply. Hence, our critique still stands.

OpenStudy (anonymous):

So... Any idea what \(T\) and \(a\) are?

OpenStudy (anonymous):

KInematics does not depend on particular planets characteristics. The laws of physics are supposed to be the same everywhere, though it is not proved, in my point of view ))).

OpenStudy (anonymous):

T is the time of flight and a is acceleration in question it is aid t=au we have to use it and find v

OpenStudy (anonymous):

You have said that \(u\) is the velocity. Is this incorrect? What does \(u\) mean?

OpenStudy (anonymous):

i mean velocity

OpenStudy (anonymous):

It is obviously a typo. Ok let's assume that t = u/a and find the formula for the velocity.

OpenStudy (anonymous):

T = total time of flight

OpenStudy (anonymous):

@ArchiePhysics

OpenStudy (anonymous):

@shivam_bhalla plzz help

OpenStudy (anonymous):

\(v=v_0+at^2\) If \(h\) is the height the particle is dropped from, we have that \(h=\frac{1}{2}at^2\) (via distance formula) Thus, \(v=v_0+2h\) \(v_0=0\) \(v=2h\) Anyone care to correct me? This seems deceptively simple.

OpenStudy (anonymous):

v can't be equal to h because of dimensions

OpenStudy (anonymous):

i will give the options 2wholeroot h/a wholeroot h/a wholeroot 2 h/a wholeroot h/2a

OpenStudy (anonymous):

@ArchiePhysics plzz just see whether u r getting by using this T=au

OpenStudy (anonymous):

\[v_f^2=v_i^2+2ad\] \[v_i=0\] \[d=h\] \[v_f^2=2ah\] \[v_f=\sqrt{2ah}\]

OpenStudy (anonymous):

There is no such option lol @Limitless

OpenStudy (anonymous):

(I'm aware it doesn't match the given answers.)

OpenStudy (anonymous):

thats i told to use T=ua

OpenStudy (anonymous):

@ArchiePhysics wat u say

OpenStudy (anonymous):

Even if admit that T = au my answer is different.

OpenStudy (anonymous):

None of those answers make sense to me. It is impossible for \(t=au\) where \(a\) and \(u\) are acceleration and velocity, respectively. Making such a statement renders our formulas nonsensical.

OpenStudy (anonymous):

Plz @Limitless u just use this a see that u get

OpenStudy (anonymous):

Just to be sure what you mean by wholeroot?

OpenStudy (anonymous):

\[\sqrt{h/a}\]

OpenStudy (anonymous):

@ArchiePhysics Like this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Maybe I am wrong, but none of these answers work out properly. (Check them dimensionally.)

OpenStudy (anonymous):

here is my solution. i haven't checked through for mistakes yet T=au assuming constant acceleration (negligible air resistance) V = U - At for upwards flight: V = 0 U= u t = T/2 0 = u - AT/2 A = 2u/T = 2/a now at a height h we have mAh gravitational potential energy, and this equals kinetic energy at bottom mAh = (1/2)mv^2 \[v = \sqrt{2Ah} = \sqrt{\frac{4h}{a}} = 2 \sqrt{\frac{h}{a}}\]

OpenStudy (anonymous):

What does \(A\) represent as compared to, say, \(a\)?

OpenStudy (anonymous):

A is the acceleration due to gravity, i just assumed a was some constant or whatever..

OpenStudy (anonymous):

T = au is not dimensionally consistent which would explain why final answers aren't if we assume something about the dimensions of a

OpenStudy (anonymous):

@eigenschmeigen you are absolutely right if a is some constant then it is easy-peasy problem!

OpenStudy (anonymous):

What do you posit that \(u\) is? And \(T\)? Openstudy said they were velocity and time, respectively.

OpenStudy (anonymous):

i assumed they were velocity and time

OpenStudy (anonymous):

Then, if \(T=au\) where \(a\) is some number, \(u\) is velocity, and \(T\) is time, we are still dimensionally inconsistent.

OpenStudy (anonymous):

a is a constant with dimension of [a] = [t/u]=[s^2/m]

OpenStudy (anonymous):

yeaah

OpenStudy (anonymous):

That makes sense, @ArchiePhysics. I am not certain, but perhaps @eigenschmeigen did solve this after all. :)

OpenStudy (anonymous):

Hey Shakir, where are you? You've got the answer. @openstudy1

OpenStudy (anonymous):

Personal curiosity: What could \(a\) be called? I don't immediately recall what the inverse of acceleration is.

OpenStudy (anonymous):

hmm interesting. i haven't encountered it as an actual quality we describe an object having. at rest or uniform motion it would be undefined

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

i didnt understand can some one help me...

OpenStudy (anonymous):

i did nt understand the last step written by @eigenschmeigen